OpenStudy (anonymous):

Two friends can paddle a canoe at 4 miles per hour in still water. One day, they went 3 miles upstream against the current. They then traveled downstream with the help of the current for the same distance. If the round trip took 2 hours, how fast was the current? Will give a medal and become a fan! Please help & please show all steps!

OpenStudy (anonymous):

@RadEn @phi

OpenStudy (phi):

what formula do you know about speed , time and distance ?

OpenStudy (anonymous):

Time= Distance/Rate ?

OpenStudy (phi):

yes, another way to write it is rate * time is distance Not sure which version we will need yet.

OpenStudy (phi):

they tell you rate is 4 miles per hour in still water

OpenStudy (phi):

I think you need common sense now... if there is a current, it will make you go faster if you go downstream and slower if you go upstream, so the rate will not be just 4 mph

OpenStudy (phi):

The question asks how fast was the current? so the speed of the current is the unknown. call the speed of the current "x".

OpenStudy (phi):

Now let's try to make an equation with this info One day, they went 3 miles upstream against the current. you know rate of the stream is x mph rate of the boat in still water is 4 mph any idea what the rate will be going upstream ?

OpenStudy (anonymous):

4-x?

OpenStudy (anonymous):

Or would it be 3/4-x instead..

OpenStudy (phi):

yes. It should make sense. if the current was 4 mph, and they went up at 4 mph, they would be standing still. or, going down stream, if you just floated (not paddling) you would still go down stream at x mph and if you go 4 mph in still water, going downstream you would go 4+x ** Or would it be 3/4-x instead.. *** I am not sure what reasoning got you that.

OpenStudy (phi):

4-x is the speed going upstream against the current

OpenStudy (anonymous):

I got the 3 because it says that they travel 3 miles upstream, I thought that had to be added into the equation somehow

OpenStudy (phi):

ok that makes sense. But I hope you agree the rate going up stream is 4-x (where x is the rate of the stream) now let's tackle they went 3 miles upstream against the current. list what you know: distance is 3, rate (upstream) is 4-x time (no idea!) and the formula Time= Distance/Rate T = D/R (easier to type) replace the letters we know with the numbers: T = 3/(4-x) ok with that ?

OpenStudy (phi):

can you do the same thing for this part downstream with the help of the current for the same distance.

OpenStudy (anonymous):

So for downstream it would just be 3/4+x then right?

OpenStudy (phi):

you should write down the whole equation: Time (downstream) = 3/(4+x) (and we should use parens otherwise it looks like ¾ +x which is different) so far we have Time (upstream)= 3/(4-x) Time(downstream) = 3/(4+x) The times will be different (I assume that is obvious... we are going slower upstream than downstream.

OpenStudy (phi):

now focus on this info the round trip took 2 hours What can you say about the sum of the time upstream plus the time downstream?

OpenStudy (anonymous):

The sum of both would be the total time it took, right? So 3/(4-x) + 3/(4+x) =4

OpenStudy (phi):

pretty close. But what does the round trip took 2 hours mean? in other words, how did you get a 4 ?

OpenStudy (anonymous):

Oh yeah I guess it would be =2 instead of =4. I guess I thought that it would be 2 hours both ways, but it says 2 round trip. Oops..

OpenStudy (phi):

2 hours for a "round trip" means 2 hours for there and back.

OpenStudy (phi):

ok, now the interpreting part of the work is done. we now switch gears into algebra mode, and solve for x in 3/(4-x) + 3/(4+x) =2

OpenStudy (anonymous):

I distribute first right?

OpenStudy (phi):

there is nothing to distribute. those parens are the denominators of fractions if you had a*(x+4) you could distribute to ax+4a but with a/(x+4) , you can't distribute. one way to simplify is multiply the whole equation (both sides and *all* terms) by (4-x) (as a first step) can you do that ?

OpenStudy (phi):

to multiply each term by (4-x) you write (4-x) next to each term

OpenStudy (anonymous):

On the top and bottom of both sides?

OpenStudy (phi):

just the tops

OpenStudy (anonymous):

3(4-x)/(4-x) + 3(4-x)/(4+x)

OpenStudy (anonymous):

Then the first numerator would be 12-3x right?

OpenStudy (phi):

also, both sides (we always do the same thing to both sides) 3(4-x)/(4-x) + 3(4-x)/(4+x) = 2(4-x) now you could distribute the 3 in the first equation, but we will do something else. there is a rule: a thing divided by itself is 1 we have (4-x) up top and a (4-x) in the bottom. that means (4-x)/(4-x) = 1 we now have 3 + 3(4-x)/(4+x) = 2(4-x) (because the (4-x)/(4-x) is 1, and 3 times 1 is just 3) next, multiply both sides (and *all terms) by (4+x) that means write (4+x) next to each term

OpenStudy (phi):

are you lost ? you should have gotten 3(4+x) +3(4-x) = 2(4-x)(4+x)

OpenStudy (phi):

now you can distribute (on the left side) and multiply out (4-x)(4+x) on the right side.

OpenStudy (anonymous):

Sorry @phi I had to do something. :/ I'm a little confused but I might be able to get it.

OpenStudy (anonymous):

Ugh math stresses me out :/

OpenStudy (anonymous):

So is it 12+3x+12+3x=2? then I combine like terms??

OpenStudy (anonymous):

Do you know if thats right @whpalmer4?

OpenStudy (whpalmer4):

Let me see: 3 + 3(4-x)/(4+x) = 2(4-x) I would multiply both sides by (4+x), as suggested: 3*(4+x) + 3(4-x)/(4+x) * (4+x) = 2(4-x)*(4+x) now we cancel out the (4+x)/(4+x) in the left side, because that equals 1: 3(4+x) + 3(4-x) = 2(4-x)(4+x) Distribute the left side: 12 + 3x + 12 - 3x = 2(4-x)(4+x) 24 = 2(4-x)(4+x) divide both sides by 2 12 = (4-x)(4+x) expand the right side 12 = 16 + 4x - 4x - x^2 12 = 16 - x^2 can you finish solving that for me?

OpenStudy (whpalmer4):

you'll get two answers, but only one of them makes sense in the context of this problem (no negative times need apply!)

OpenStudy (whpalmer4):

we're within sight of the goal, don't up now! :-)

OpenStudy (anonymous):

I got to -28=-x^2 but I'm not sure where to go from there

OpenStudy (anonymous):

Is that right @whpalmer4 ?

OpenStudy (whpalmer4):

12 = 16 - x^2 let's subtract 12 from both sides

OpenStudy (anonymous):

14-x^2

OpenStudy (whpalmer4):

no... 12 = 16 - x^2 subtract 12 from both sides 12 - 12 = 16 - x^2 - 12 can you simplify that for me?

OpenStudy (anonymous):

12=16-x^2 12=-x^2+16 (2, -2)

OpenStudy (whpalmer4):

right, x = 2 and x = -2 are the two solutions to that equation. which one makes sense for this problem?

OpenStudy (anonymous):

Um x=2 I think

OpenStudy (anonymous):

Well yeah it would have to be positive 2.

OpenStudy (whpalmer4):

right. as a review of the early pages of this discussion shows, x was the speed of the stream in miles per hour. so, going with the current, we go 2+4=6mph, and against we go 4-2 = 2 mph. now, does that answer satisfy the problem?

OpenStudy (anonymous):

So this means that the current was going 2 mph?

OpenStudy (whpalmer4):

let's check. 3 miles with the current, and 3 miles against the current, and the total time was 2 hours. with the current, to cover 3 miles @ 6 mph = 3 miles / 6 (miles/hr) = 1/2 hour against the current, to cover 3 miles @ 2 mph = 3 miles / 2 (miles/hr) = 1 1/2 hours total time = 1/2 hour + 1 1/2 hour = 2 hours our answer checks out!

OpenStudy (whpalmer4):

yes, x was the speed of the current.

OpenStudy (anonymous):

Thanks so much for your help :))

OpenStudy (whpalmer4):

you're welcome! I like these problems :-) @phi did a good job setting up the problem with you, too.

OpenStudy (anonymous):

Yes, thank you @phi!! He's helped me with a few before :) You are both great!